# Learning Mamba as a Continual Learner

### v1

@@ 3.1 Preliminaries: Transformers, Linear Transformers, and SSMs
Transformers produce next-token predictions in sequence relying on a self-attention mechanism (Vaswani et al., 2017). Given a sequence of $N$ vectors in $M$-dimension denoted as ${\mathbf{Z}}\in{\mathbb{R}}^{N\times M}$, the vanilla self-attention is formulate with a softmax attention method:
${\mathbf{Q}}\!={\mathbf{Z}}{\mathbf{W}}_{Q},~{}{\mathbf{K}}\!={\mathbf{Z}}{\mathbf{W}}_{K},~{}{\mathbf{V}}\!={\mathbf{Z}}{\mathbf{W}}_{V},~{}~{}{\mathbf{u}}_{t}=\sum_{j=1}^{N}\ \frac{\exp\left({\mathbf{Q}}_{t}{\mathbf{K}}_{j}^{\top}/{\sqrt{d}}\right)}{\sum_{j=1}^{N}\ \exp\left({\mathbf{Q}}_{t}{\mathbf{K}}_{j}^{\top}/{\sqrt{d}}\right)}{\mathbf{V}}_{j},$ (1)
@@
where ${\mathbf{W}}_{Q}\in{\mathbb{R}}^{M\times C},~{}{\mathbf{W}}_{K}\in{\mathbb{R}}^{M\times C},~{}{\mathbf{W}}_{V}\in{\mathbb{R}}^{C\times C}$ are the projection weight matrices, ${\mathbf{u}}_{t}\in{\mathbb{R}}^{C}$ denote the output embedding, and $C$ is the hidden dimension. ${\mathbf{Q}}_{t}$, ${\mathbf{K}}_{j}$, and ${\mathbf{V}}_{j}$ denote the indexed vectors in the corresponding matrices. The notation fonts are slightly abused to be consistent with the literatures. Each input token generates a key-value pair, leading to a linearly increased key-value cache size. Softmax attention measures the similarities between the query-key pairs, leading to ${\mathcal{O}}(N^{2})$ complexity.
@@
Linear Transformer (Katharopoulos et al., 2020) reduces the complexity relying a linear attention method. By applying a feature representation function $\phi()$ corresponding to a kernel for ${\mathbf{Q}}$ and ${\mathbf{K}}$, the linear attention method replaces the softmax attention with a linear operation as:
${\mathbf{u}}_{t}=\sum_{j=1}^{N}\ \frac{{\mathbf{Q}}_{t}{\mathbf{K}}_{j}^{\top}}{\sum_{j=1}^{N}\!{{\mathbf{Q}}_{t}{\mathbf{K}}_{j}^{\top}}}{\mathbf{V}}_{j}=\frac{{\mathbf{Q}}_{t}\!\left(\sum_{j=1}^{N}\!{{\mathbf{K}}_{j}^{\top}{\mathbf{V}}_{j}}\right)}{{\mathbf{Q}}_{t}\!\left(\ \sum_{j=1}^{N}\!{{\mathbf{K}}_{j}^{\top}}\right)},$ (2)
where ${\mathbf{Q}}\!=\!\phi({\mathbf{Z}}{\mathbf{W}}_{Q}),~{}{\mathbf{K}}\!=\!\phi({\mathbf{Z}}{\mathbf{W}}_{K}),~{}{\mathbf{V}}\!\!={\mathbf{Z}}{\mathbf{W}}_{V}$, $\phi(\cdot)$ is set as $\phi({\mathbf{x}})=\mathrm{elu}({\mathbf{x}})+1$ in (Katharopoulos et al., 2020). Performer employs $\phi({\mathbf{x}})=\exp\left({\mathbf{x}}{\mathbf{W}}_{p}-\|{\mathbf{x}}\|^{2}/2\right)$, with ${\mathbf{W}}_{p}$ comprising orthogonal random vectors (Choromanski et al., 2020). Through rearranging $({\mathbf{Q}}{\mathbf{K}}^{\top}){\mathbf{V}}$ as ${\mathbf{Q}}({\mathbf{K}}^{\top}{\mathbf{V}})$ according to associative property, the computational complexity is reduced to ${\mathcal{O}}(N)$.
@@
In practice, the attention operations in Eq. (1) and (2) can be implemented in autoregressive models, where calculation of ${\mathbf{u}}_{t}$ can only see the proceeding tokens with $j\leq t$. Specifically, with the causal masking, the linear attention can be rewritten as:
${\mathbf{u}}_{t}=\frac{{\mathbf{Q}}_{t}\left(\sum_{j=1}^{t}{{\mathbf{K}}_{j}^{\top}{\mathbf{V}}_{j}}\right)}{{\mathbf{Q}}_{t}\left(\ \sum_{j=1}^{t}{{\mathbf{K}}_{j}^{\top}}\right)}=\frac{{\mathbf{Q}}_{t}{\mathbf{S}}_{t}}{{\mathbf{Q}}_{t}{\mathbf{G}}_{t}},~{}~{}~{}{\mathbf{S}}_{t}={\mathbf{S}}_{t-1}+{{\mathbf{K}}_{t}^{\top}{\mathbf{V}}_{t}},~{}{\mathbf{G}}_{t}={\mathbf{G}}_{t-1}+{{\mathbf{K}}_{t}^{\top}},$ (3)
@@
In the autoregressive process with causal masking, the softmax attention operation Eq. (1) in Transformer can be seen as a recurrent process based on an accumulated set of key-value pairs $\{({\mathbf{K}}_{j},{\mathbf{V}}_{j})\}_{j=1}^{t}$ as a hidden state (Katharopoulos et al., 2020).
Structured state space sequence models (SSM or S4) (Gu & Dao, 2023; Gu et al., 2021a; Dao & Gu, 2024) are sequence models describing a system that maps input $z_{t}\in{\mathbb{R}}$ to output $u_{t}\in{\mathbb{R}}$ through a hidden state ${\mathbf{h}}_{t}\in{\mathbb{R}}^{C\times 1}$ in a discrete sequence applied with neural networks. Specifically, SSMs can be formulated with parameters ${\mathbf{A}}\in{\mathbb{R}}^{C\times C}$, ${\mathbf{B}}\in{\mathbb{R}}^{C\times 1}$, ${\mathbf{C}}\in{\mathbb{R}}^{1\times C}$, and $D\in{\mathbb{R}}$, as
${\mathbf{h}}_{t}={\mathbf{A}}{\mathbf{h}}_{t-1}+{\mathbf{B}}z_{t},~{}~{}u_{t}={\mathbf{C}}{\mathbf{h}}_{t}+Dz_{t}.$ (4)
We directly formulate a discrete SSM in Eq. (4), where the ${\mathbf{A}}$ and ${\mathbf{B}}$ are transformed from a continuous version ${\mathbf{A}}^{\prime}$ and ${\mathbf{B}}^{\prime}$ relying on a timescale parameter $\Delta\in{\mathbb{R}}$, via ${\mathbf{A}}=\exp(\Delta{\mathbf{A}}^{\prime})$ and ${\mathbf{B}}={\mathbf{A}}^{-1}({{\mathbf{A}}}-{\mathbf{I}})\cdot\Delta{\mathbf{B}}^{\prime}$. ${\mathbf{A}}$ and ${\mathbf{B}}$ perform the selection or gating in hidden state updating.
Figure 2: Illustration of the designs of Mamba block.
@@ 3.2 SSM and Mamba for Meta-Continual Learning
Selective SSM & Mamba in MCL. The dynamics of the basic SSM or S4 are time-invariant, restricting the model’s ability to handle complex sequences. Mamba (Gu & Dao, 2023) incorporates a selective SSM into the model by generating the input-dependent SSM parameters to reflect the input/step-sensitive selection process. The selective SSM can be written as:
@@
where ${\mathbf{A}}_{t}$, ${\mathbf{B}}_{t}$, and ${\mathbf{C}}_{t}$ are produced in Mamba relying on the input token at step $t$. Different from Transformers maintaining key-value pairs for all input tokens (leading to linearly increasing state size), Mamba compresses the context information in a fixed/constant-size hidden state, matching the efficiency requirements and original objective of CL.
In our MCL tasks and other practical scenarios, we need Mamba to handle the input sequence ${\mathbf{Z}}\in{\mathbb{R}}^{N\times M}$ with each token as a vector ${\mathbf{z}}_{t}\in{\mathbb{R}}^{M}$. Mamba applies the selective SSM to each dimension/channel independently:
${\mathbf{H}}_{t}=[{\mathbf{A}}_{t,i}{\mathbf{h}}_{t-1,i}+{\mathbf{B}}_{t,i}{\mathbf{z}}_{t,i}]_{i=1}^{M},~{}~{}{\mathbf{u}}_{t}={\mathbf{C}}_{t}{\mathbf{H}}_{t}+{\mathbf{D}}\odot{\mathbf{z}}_{t},$ (6)
where ${\mathbf{H}}_{t}\in{\mathbb{R}}^{C\times M}$ is a concatenation of the hidden state corresponding to all $M$ dimensions of the input embedding, ${\mathbf{C}}_{t}\in{\mathbb{R}}^{1\times C}$, ${\mathbf{D}}\in{\mathbb{R}}^{1\times M}$, and ${\mathbf{u}}_{t}\in{\mathbb{R}}^{1\times M}$. As shown in Fig. 2, the Mamba block used in our work applies a 1-D convolution on the input tokens and then projects the representations to obtain the input-dependent SSM parameters (Dao & Gu, 2024). Multiple Mamba blocks are stacked homogeneously. Relying on the selective mechanism (Gu & Dao, 2023), Mamba’s ability to handle complex MCL tasks can be stronger than other attention-free models and competitive or better than Transformers with a key-value cache.
@@ 3.2.1 MCL with Mamba as a Continual Learner
We will train a Mamba model $f_{\theta}()$ to perform CL by processing an online data stream containing paired $({\mathbf{x}},y)$; the model can produce predictions relying on the retained hidden state, for all the seen tasks. Meta-learning of such continual learner will be conducted on multiple CL episodes. Each CL episode contains a training data stream ${\mathcal{D}}^{\text{train}}$ and a testing set ${\mathcal{D}}^{\text{test}}$ from the same task distribution, denoted as $P_{({\mathcal{X}},{\mathcal{Y}})}$ with $({\mathcal{D}}^{\text{train}},{\mathcal{D}}^{\text{test}})\sim P_{({\mathcal{X}},{\mathcal{Y}})}$. For example, in ICL, all classes used for testing should have been seen in the preceding classes in the data stream. The objective of MambaCL is to meta-learn the parameter of Mamba model, i.e., $\theta$, to perform prediction $\hat{y}^{\text{test}}=f_{\theta}(({\mathcal{D}}^{\text{train}},{\mathbf{x}}^{\text{test}}))$ for any $({\mathbf{x}}^{\text{test}}_{i},{\mathbf{y}}^{\text{test}}_{i})\in{\mathcal{D}}^{\text{test}}$. The CL task can be treated as next token prediction problem in sequence: $({\mathbf{x}}^{\text{train}}_{1},{\mathbf{y}}^{\text{train}}_{1},...,{\mathbf{x}}^{\text{train}}_{T},{\mathbf{y}}^{\text{train}}_{T},{\mathbf{x}}^{\text{test}}_{k})\rightarrow y^{\text{test}}_{k}$. The meta-learning of a Mamba continual learner can be performed by optimizing the sequence prediction task on a series of sampled CL episodes:
@@
On the data stream, the meta-learned Mamba $f_{\theta}()$ recognizes the association relationship between ${\mathbf{x}}$ and $y$ through the sequence, and then recurrently updates the hidden state ${\mathbf{H}}_{t}$, which can used for prediction, as shown in Fig. 1. This efficient online CL process selects and compresses the knowledge in the data stream in a time-variant and content-aware selective manner. To further validate the extension ability of Mamba in MCL, we also explore the potential of incorporating mixture-of-expert (MoE) architecture into Mamba model (Fedus et al., 2022; Pioro et al., 2024) for learning and mixing multiple learners.
@@
Target token embeddings. The value of the target $y$ is essentially a symbol with consistent indication meaning for ${\mathbf{x}}$ within each episode, which does not take any global meaning across the episode. The model is thus trained to handle arbitrary CL episodes with the ability to generalize to different domains. Instead of pre-defining a small and fixed feasible set of candidate targets e.g., classes, and a restricted prediction head, we conduct token embeddings for targets based on a universal and large vocabulary (Lee et al., 2023), inspired by the tokenization in LMs (Sennrich, 2015; Devlin et al., 2019). For each episode, a subset of unique codes is randomly picked from the vocabulary to indicate different classes; in inference, the sequence model produces the probability of the next step for all possible tokens in the vocabulary. Instead of conducting experiments of meta-training and meta-testing with the same number of classes (Lee et al., 2023), we conduct generalization analyses.
3.2.2 Regularizing Selectivity of Mamba for Meta-Training

### v4

@@ 3.1 Preliminaries: Transformers and SSMs
Transformers produce next-token predictions in sequence relying on a self-attention mechanism [65]. Given a sequence of $N$ vectors in $M$-dimension denoted as ${\mathbf{Z}}\in{\mathbb{R}}^{N\times M}$, the vanilla self-attention is formulate with a softmax attention method:
$\begin{split}{\mathbf{Q}}\!={\mathbf{Z}}{\mathbf{W}}_{Q},~{}{\mathbf{K}}\!={\mathbf{Z}}{\mathbf{W}}_{K},~{}{\mathbf{V}}\!={\mathbf{Z}}{\mathbf{W}}_{V},~{}~{}~{}{\mathbf{u}}_{t}=\sum_{j=1}^{N}\ \frac{\exp\left({\mathbf{Q}}_{t}{\mathbf{K}}_{j}^{\top}/{\sqrt{d}}\right)}{\sum_{j=1}^{N}\ \exp\left({\mathbf{Q}}_{t}{\mathbf{K}}_{j}^{\top}/{\sqrt{d}}\right)}{\mathbf{V}}_{j},\end{split}$ (1)
@@
where ${\mathbf{W}}_{Q}\in{\mathbb{R}}^{M\times C},~{}{\mathbf{W}}_{K}\in{\mathbb{R}}^{M\times C},~{}{\mathbf{W}}_{V}\in{\mathbb{R}}^{M\times M}$ are the projection weight matrices, ${\mathbf{u}}_{t}\in{\mathbb{R}}^{M}$ denote the output embedding. ${\mathbf{Q}}_{t}$, ${\mathbf{K}}_{j}$, and ${\mathbf{V}}_{j}$ denote the indexed vectors in the corresponding matrices. The notation fonts are slightly abused for consistency with conventions. Each input token generates a key-value pair, leading to a linear increase in key-value cache size. Softmax attention measures the similarities between the query-key pairs, leading to ${\mathcal{O}}(N^{2})$ complexity.
@@
Linear Transformer [27] reduces the complexity relying on a linear attention method. By applying a feature representation function $\phi()$ corresponding to a kernel for ${\mathbf{Q}}$ and ${\mathbf{K}}$, the linear attention method replaces the softmax attention with a linear operation as:${\mathbf{u}}_{t}=\sum_{j=1}^{N}\frac{{\mathbf{Q}}_{t}{\mathbf{K}}_{j}^{\top}}{\sum_{j^{\prime}=1}^{N}{\mathbf{Q}}_{t}{\mathbf{K}}_{j^{\prime}}^{\top}}{\mathbf{V}}_{j}=\frac{{\mathbf{Q}}_{t}\left(\sum_{j=1}^{N}{\mathbf{K}}_{j}^{\top}{\mathbf{V}}_{j}\right)}{{\mathbf{Q}}_{t}\left(\sum_{j=1}^{N}{\mathbf{K}}_{j}^{\top}\right)}$, where ${\mathbf{Q}}\!=\!\phi({\mathbf{Z}}{\mathbf{W}}_{Q}),~{}{\mathbf{K}}\!=\!\phi({\mathbf{Z}}{\mathbf{W}}_{K}),~{}{\mathbf{V}}\!\!={\mathbf{Z}}{\mathbf{W}}_{V}$, $\phi(\cdot)$ is set as $\phi({\mathbf{x}})=\mathrm{elu}({\mathbf{x}})+1$ in [27]. Performer employs $\phi({\mathbf{x}})=\exp\left({\mathbf{x}}{\mathbf{W}}_{p}-\|{\mathbf{x}}\|^{2}/2\right)$, with ${\mathbf{W}}_{p}$ comprising orthogonal random vectors [8]. The computational complexity is reduced to ${\mathcal{O}}(N)$.
@@
In practice, the attenstion operations can be implemented in autoregressive models, where calculation of ${\mathbf{u}}_{t}$ can only see the proceeding tokens with $j\leq t$. Specifically, with the causal masking, the linear attention can be rewritten as:
$\begin{split}{\mathbf{u}}_{t}=\frac{{\mathbf{Q}}_{t}\left(\sum_{j=1}^{t}{{\mathbf{K}}_{j}^{\top}{\mathbf{V}}_{j}}\right)}{{\mathbf{Q}}_{t}\left(\ \sum_{j=1}^{t}{{\mathbf{K}}_{j}^{\top}}\right)}=\frac{{\mathbf{Q}}_{t}{\mathbf{S}}_{t}}{{\mathbf{Q}}_{t}{\mathbf{G}}_{t}},~{}~{}~{}{\mathbf{S}}_{t}={\mathbf{S}}_{t-1}+{{\mathbf{K}}_{t}^{\top}{\mathbf{V}}_{t}},~{}{\mathbf{G}}_{t}={\mathbf{G}}_{t-1}+{{\mathbf{K}}_{t}^{\top}},\end{split}$ (2)
@@
State space sequence models (SSMs) [16, 17, 9] are sequence models describing a system that maps input $z_{t}\in{\mathbb{R}}$ to output $u_{t}\in{\mathbb{R}}$ through a hidden state ${\mathbf{h}}_{t}\in{\mathbb{R}}^{C\times 1}$ in a discrete sequence applied with neural networks. Specifically, SSMs can be formulated with parameters ${\mathbf{A}}\in{\mathbb{R}}^{C\times C}$, ${\mathbf{B}}\in{\mathbb{R}}^{C\times 1}$, ${\mathbf{C}}\in{\mathbb{R}}^{1\times C}$, and $D\in{\mathbb{R}}$, as ${\mathbf{h}}_{t}={\mathbf{A}}{\mathbf{h}}_{t-1}+{\mathbf{B}}z_{t},\quad u_{t}={\mathbf{C}}{\mathbf{h}}_{t}+Dz_{t}$. We can directly formulate a discrete SSM, where the ${\mathbf{A}}$ and ${\mathbf{B}}$ are transformed from a continuous version ${\mathbf{A}}^{\prime}$ and ${\mathbf{B}}^{\prime}$ relying on a timescale parameter $\Delta\in{\mathbb{R}}$, via ${\mathbf{A}}=\exp(\Delta{\mathbf{A}}^{\prime})$ and ${\mathbf{B}}={\mathbf{A}}^{-1}({{\mathbf{A}}}-{\mathbf{I}})\cdot\Delta{\mathbf{B}}^{\prime}$. ${\mathbf{A}}$ and ${\mathbf{B}}$ perform the selection or gating in hidden state updating.
@@ 3.2 SSM and Mamba for Meta-Continual Learning
Selective SSM & Mamba in MCL. The dynamics of the basic SSM are time-invariant, limiting their ability to model complex sequences. Mamba [16] incorporates a selective SSM into the model by generating input-dependent SSM parameters to reflect the input/step-sensitive selection process. The selective SSM can be written as:
@@
where ${\mathbf{A}}_{t}$, ${\mathbf{B}}_{t}$, and ${\mathbf{C}}_{t}$ are produced in Mamba relying on the input token at step $t$. Unlike Transformers storing key-value pairs for all input tokens (causing linear state growth), Mamba compresses context into a fixed-size hidden state, aligning with CL’s efficiency goals.
In our MCL tasks and other practical scenarios, we need Mamba to handle the input sequence ${\mathbf{Z}}\in{\mathbb{R}}^{N\times M}$ with each token as a vector ${\mathbf{z}}_{t}\in{\mathbb{R}}^{M}$. Mamba applies the selective SSM to each dimension/channel independently with:
${\mathbf{H}}_{t}=[{\mathbf{A}}_{t,i}{\mathbf{h}}_{t-1,i}+{\mathbf{B}}_{t,i}{\mathbf{z}}_{t,i}]_{i=1}^{M},~{}{\mathbf{u}}_{t}={\mathbf{C}}_{t}{\mathbf{H}}_{t}+{\mathbf{D}}\odot{\mathbf{z}}_{t},$ (4)
where ${\mathbf{H}}_{t}\in{\mathbb{R}}^{C\times M}$ is the concatenated hidden state across all $M$ input dimensions, ${\mathbf{C}}_{t}\in{\mathbb{R}}^{1\times C}$, ${\mathbf{D}}\in{\mathbb{R}}^{1\times M}$, and ${\mathbf{u}}_{t}\in{\mathbb{R}}^{1\times M}$. As shown in Fig. 2, the Mamba block used in our work applies a 1-D convolution on the input tokens and then projects the representations to obtain input-dependent SSM parameters [9]. Multiple Mamba blocks are stacked homogeneously. With its selective mechanism [16], Mamba demonstrates strong performance in complex MCL tasks, outperforming other attention-free models and showing competitive results compared to Transformers with key-value caching.
@@ 3.2.1 MCL with Mamba as a Continual Learner
We meta-learn Mamba model $f_{\theta}()$ as a continual learner using multiple CL streaming epidodes, where the models learning and prediction as SP relying the retained hidden state within each episode. Each CL episode contains a training data stream ${\mathcal{D}}^{\text{train}}$ and a testing set ${\mathcal{D}}^{\text{test}}$ from the same task distribution, denoted as $P_{({\mathcal{X}},{\mathcal{Y}})}$ with $({\mathcal{D}}^{\text{train}},{\mathcal{D}}^{\text{test}})\sim P_{({\mathcal{X}},{\mathcal{Y}})}$. For instance, in ICL, all testing classes have been seen in the preceding data stream. The objective of MambaCL is to meta-learn the Mamba model’s parameter, i.e., $\theta$, to perform prediction $\hat{y}^{\text{test}}=f_{\theta}(({\mathcal{D}}^{\text{train}},{\mathbf{x}}^{\text{test}}))$ for any $({\mathbf{x}}^{\text{test}}_{i},{\mathbf{y}}^{\text{test}}_{i})\in{\mathcal{D}}^{\text{test}}$. The CL task can be framed as the next token prediction problem in sequence: $({\mathbf{x}}^{\text{train}}_{1},{\mathbf{y}}^{\text{train}}_{1},...,{\mathbf{x}}^{\text{train}}_{T},{\mathbf{y}}^{\text{train}}_{T},{\mathbf{x}}^{\text{test}}_{k})\rightarrow y^{\text{test}}_{k}$. Meta-learning a Mamba continual learner involves optimizing this SP task across multiple sampled CL episodes:
@@
On the data stream, the meta-learned Mamba $f_{\theta}()$ recognizes the association relationship between ${\mathbf{x}}$ and $y$ through the sequence, recurrently updating the hidden state ${\mathbf{H}}_{t}$ for prediction, as shown in Fig. 1. This efficient online CL process selects and compresses the knowledge in the data stream in a time-variant and content-aware selective manner. To further validate the extension ability of Mamba in MCL, we also explore the potential of incorporating mixture-of-experts (MoE) into the Mamba model [13, 47] for learning and mixing multiple learners (See Sec. 4.3).
@@
Target token embeddings. The value of the target $y$ is essentially a symbol with consistent indication meaning for ${\mathbf{x}}$ within each episode, which does not take any global meaning across the episode. The model is thus trained to handle arbitrary CL episodes with the ability to generalize to different domains. Instead of pre-defining a small and fixed feasible set of candidate targets e.g., classes, and a restricted prediction head, we conduct token embeddings for targets based on a universal and large vocabulary [32], inspired by the tokenization in LMs [54, 12]. For each episode, a subset of unique codes is randomly picked from the vocabulary to indicate different classes; in inference, the sequence model produces the probability of the next step for all possible tokens in the vocabulary. Beyond conducting experiments of meta-training and meta-testing with the same number of classes [32], we conduct generalization analyses in Sec. 4.2.
3.2.2 Selectivity Regularization for Meta Training
Meta-learning a continual learner to associate inputs and targets from a data stream is challenging for both attention-free models and Transformers [32]. Meta-training can be slow and challenging to converge due to long-range dependencies in supervision on the stream. We thus provide additional guidance by strengthening associations (known in the meta-training dataset) between query tokens (i.e., testing inputs) and their correlated preceding tokens. While straightforward for Transformers with explicit intra-sequence attention [32], this process is non-trivial for SSM-based Mamba.